find the two consecutive even integers whose product is 624

Answer by mananth(16949)

let x be the intgere
the next even number = x+2
x(x+2)= 624
x^2+2x-624=0
x^2+26x-24x-424=0
x(x+26)-24(x+26)=0
(x+26)(x-24)=0
x= -26 OR 24
The numbers are
-26,&-24
24 & 26